OpenStudy (anonymous):

Given the function f(x) = 4(x + 1)^2 − 3, indicate the shifts that will affect the location of the vertex and explain what effect they will have. Use complete sentences. f(x−2) f(x) − 2 f(2x) 2•f(x)

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@amistre64

OpenStudy (math&ing001):

f(x-2)=4((x-2)+1)^2-3=4(x-1)^2-3. So basicly you just replace the x with the value between brackets.

OpenStudy (anonymous):

Okay, I was understanding that but, I couldn't figure out what the equation would be

OpenStudy (math&ing001):

Which equation are talking about ?

OpenStudy (anonymous):

f(2x)

OpenStudy (math&ing001):

That would be f(2x)=4(2x+1)^2-3

OpenStudy (anonymous):

Okay, I see how it works. Thank you!

OpenStudy (math&ing001):

you're welcome !

OpenStudy (anonymous):

f(x-2)=4((x-2)+1)^2-3=4(x-1)^2-3

OpenStudy (anonymous):

In that equation, why is there two "=4's"